OpenStudy (anonymous):

What polynomial has roots of -5, 2, and 4 ? I don't know how to do this at all.

OpenStudy (retireed):

Think about factoring and roots. (x + Y) = 0 this is how you find roots for this to work X must equal -Y

OpenStudy (retireed):

Now use the numbers given in the question and substitute each in for X and then figure out what Y must be. Remember they gave you the roots and you want to find the (x+y) or (x-z).

OpenStudy (anonymous):

Okay if I plug the numbers in (-5+y) (2+y) (4+y) would I make then equal zero?

OpenStudy (retireed):

You have three roots so you need three (x +y) factors They might want to multiply the three factors to come up with the polynomial.... ax^3 + bx^2 + cx + d = 0

OpenStudy (anonymous):

ok so -5x^3+2x^2+4xx+d=0

OpenStudy (retireed):

Yes (to answer your question) (-5+y) = 0 y = ?? (2+y) = 0 y = ?? (4+y) = 0 y = ??

OpenStudy (anonymous):

y=5 y=-2 y=-4

OpenStudy (retireed):

Yes and now ... (x + 5) ( x -2) (x -4) multiply these

OpenStudy (anonymous):

is it x^3+40?

OpenStudy (retireed):

No it will look like this ax^3 + bx^2 + cx + d = 0 do it parts (x + 5) ( x -2) = quad then multiply quad and (x -4)

OpenStudy (retireed):

Do the (x + 5) ( x -2) ??

OpenStudy (anonymous):

I got fro that one x^2+3x-10

OpenStudy (anonymous):

How would I do the other one?

OpenStudy (retireed):

x^2 + 3x - 10 that is what i got

OpenStudy (retireed):

Now just multiply (x^2 + 3x - 10) (x -4) one term at a time by that I mean x^2 times x then 3x times x three -10 times x then do the same thing only multiply by -4 simplify by collecting the terms

OpenStudy (anonymous):

x^2+3x^2+40

OpenStudy (retireed):

No x^2 times x = x^2 * x^1 = x^? add the exponents

OpenStudy (retireed):

or look at as x* x* x = x^? remember like with real numbers 2 * 2 * 2 = 8 = 2^3

OpenStudy (anonymous):

so it would be x^3+3x^2_____+40???

OpenStudy (anonymous):

These are my answer choices Select one: A. x^3 - x^2 - 22x + 40 B. x^3 + x^2 - 22x - 40 C. x^3 + 3x^2 - 18x - 40 D. x^3 - 3x^2 - 18x + 40

OpenStudy (retireed):

One more time you need to multiply all three terms by the -4 also

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

-18x

OpenStudy (retireed):

I need to draw it out so you understand this. I want you to understand the process. I will try to be quick.

OpenStudy (retireed):